OpenStudy (anonymous):

What is the solution to the equation? help please? √(2x+10)-6=2

OpenStudy (campbell_st):

add 6 to both sides then square both sides to remove the radical then solve for x in the normal way

OpenStudy (anonymous):

so by doing that i get 2x+10=8? then what do i di?

OpenStudy (campbell_st):

not quite you need to have 2x + 10 = 8^2

OpenStudy (anonymous):

x=27??